Crew 210 Sol Summary Report 17-APR-2019
Summary Title: Plants to the north and plants to the west
Author’s name: Paul Sokoloff
Mission Status: Nominal
Sol Activity Summary: Today we continued our outreach efforts on social media, including answering questions on our Reddit "Ask a Martian Anything", and uploaded a Facebook update video. Four of the crew went to Copernicus Valley to continue transect and quadrat sampling, where we discovered new plant records for the MDRS area, including a stinkweed (Cleomella palmeriana). After a beautiful ride out and back, we returned home to the hab for lunch. In the afternoon we hiked up to the top of Hab Ridge just to the west of MDRS to finish our quadrat sampling. We returned to the hab around 6, and we plan on spending the evening processing samples and pressing plants in the Science Dome.
Look Ahead Plan: Tomorrow AM we will conduct livestream outreach broadcasts (one in French and one in English) in partnership with Discover The Universe and Live It. In the afternoon we will continue to process samples in the science dome, and will meet to discuss next steps for our multi-year program.
Anomalies in work: No problems.
Weather: Bright and Sunny
Crew Physical Status: Good. One crew member slipped into a sinkhole on Sagan Street but caught themselves before falling far, and is OK (one very slight scrape). No problems afterwards
EVA: Four crew members took ATVs to Copernicus Valley in the northwest corner of the MDRS operation area in the morning for quadrat sampling. In the afternoon four crew members hiked up to Hab Ridge for additional quadrat surveys.
Reports to be file: GreenHab Report, Operations Report, Photo Report
Support Requested: None.